# Lifting the Veil on the Large Language Model Supply Chain: Composition, Risks, and Mitigations

### v1

@@ Abstract.
Large language models (LLM) have sparked significant impact with regard to both intelligence and productivity. In recent years, a great surge has been witnessed in the introduction of both commercial and open-source LLMs. Many businesses have adopted the LLMs into their applications to solve their own domain-specific tasks. However, integrating LLMs into specific business scenarios requires more than just utilizing the models themselves. Instead, it is a systematic process that involves substantial components, which are collectively referred to as the LLM supply chain. The LLM supply chain inherently carries risks. Therefore, it is essential to understand the types of components that may be introduced into the supply chain and the associated risks, enabling different stakeholders to implement effective mitigation measures. While some literature touches on risks associated with the LLM supply chain, there is currently no paper that explicitly defines its scope, identifies inherent risks, and examines potential mitigation strategies. As LLMs have become essential infrastructure in the new era, we believe that a thorough review of the LLM supply chain, along with its inherent risks and mitigation strategies, would be valuable for industry practitioners to avoid potential damages and losses, and enlightening for academic researchers to rethink existing approaches and explore new avenues of research. Our paper provides a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types. We developed taxonomies of risk types, risky actions, and mitigations related to various supply chain stakeholders and components. In summary, our work explores the technical and operational aspects of the LLM supply chain, offering valuable insights for researchers and engineers in the evolving LLM landscape.
@@ 1. Introduction
While some literature touches on risks associated with the LLM supply chain, there is currently no paper that explicitly defines its scope, identifies inherent risks, and examines potential mitigation strategies in the LLM supply chain. Ladisa et al. (Ladisa et al., 2023) developed taxonomies for attack and defense techniques in software supply chains. However, their taxonomies do not account for LLM-specific components such as pre-trained LLMs, prompts, and data, making their survey less attractive for LLM supply chain stakeholders. Several surveys address the security and privacy challenges (Yao et al., 2024a; Cui et al., 2024b; Kumar et al., 2024; Yao et al., 2024a; Neel and Chang, 2023; Esmradi et al., 2023) and discuss the defense and mitigation strategies of LLMs. However, they did not discuss further what kinds of stakeholders and artifacts are affected, which fails to provide actionable insights and a clear scope to the corresponding stakeholders. Additionally, Wang et al. first (Wang et al., 2024b) outline a future research agenda for the LLM supply chain. However, they did not clearly define the scope of the elements that make up the LLM supply chain, making the covered challenges and opportunities that are less comprehensive than they could be. As the topic is still emerging and rapidly evolving, existing works may not comprehensively incorporate insightful ideas and perspectives from online preprints or web blogs. Therefore, we believe that a comprehensive LLM supply chain overview as well as a taxonomy of risks and clear guidance of mitigation in the LLM supply chain could be beneficial for both industrial and academic audiences. It could guide practitioners to pay attention to the LLM supply chain risks and provide a visionary map for academic researchers to devise new mitigation or defensive techniques.
In our paper, we provide a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types, as illustrated in Figure 1. We developed taxonomies of risk types, risky actions, and mitigations related to various LLM supply chain stakeholders and components. Specifically, we fist holistically collected both academic and online resources. Then, we summarize the risk scenarios as stakeholders performing risky actions on specific supply chain components, which in turn lead to distinct risk types. We illustrate the risk scenarios from the perspective of upstream contributors, downstream users, and administrators. Furthermore, we list the types of various mitigation measures in response to the LLM supply chain risks.
@@ 2.2.5. Applications
Applications provide functionalities and services directly used by users. These applications include chatbots, automated customer service, text-generation tools, and more. Application developers integrate large models and API interfaces into real-world products to develop intelligent applications that meet specific needs. These applications directly serve end users, improving work efficiency or solving practical problems.
@@ 4.3. Taxonomy of Risky Actions
4.3.1. Risks on Third-party Libraries
Third-party libraries, including packages, frameworks, and plugins, have become essential at every stage of LLM application development. However, using these libraries comes with potential risks.
Implant Malicious Code. Attackers can develop malicious code and attempt to implant it through various methods, finally integrating it into the LLM supply chain. It can cause the risk of malicious use (LR2).

### v2

@@ Abstract.
Large language models (LLM) have sparked significant impact with regard to both intelligence and productivity. In recent years, a great surge has been witnessed in the introduction of both commercial and open-source LLMs. Many businesses have adopted the LLMs into their applications to solve their own domain-specific tasks. However, integrating LLMs into specific business scenarios requires more than just utilizing the models themselves. Instead, it is a systematic process that involves substantial components, which are collectively referred to as the LLM supply chain. The LLM supply chain inherently carries risks. Therefore, it is essential to understand the types of components that may be introduced into the supply chain and the associated risks, enabling different stakeholders to implement effective mitigation measures. While some literature discusses risks associated with LLMs, there is currently no paper that clearly outlines the LLM supply chain from the perspective of both providing and consuming its components. As LLMs have become essential infrastructure in the new era, we believe that a thorough review of the LLM supply chain, along with its inherent risks and mitigation strategies, would be valuable for industry practitioners to avoid potential damages and losses, and enlightening for academic researchers to rethink existing approaches and explore new avenues of research. Our paper provides a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types. We developed taxonomies of risk types, risky actions, and mitigations related to various supply chain stakeholders and components. In summary, our work explores the technical and operational aspects of the LLM supply chain, offering valuable insights for researchers and engineers in the evolving LLM landscape.
@@ 1. Introduction
While some literature discusses risks associated with LLMs, there is currently no paper that clearly outlines the LLM supply chain from the perspective of both providing and consuming its components. As a result, the inherent risks and potential mitigation strategies are less organized and comprehensive. Ladisa et al. (Ladisa et al., 2023) developed taxonomies for attack and defense techniques in software supply chains. However, their taxonomies do not account for LLM-specific components such as pre-trained LLMs, prompts, and data, making their survey less attractive for LLM supply chain stakeholders. Several surveys address the security and privacy challenges (Yao et al., 2024a; Cui et al., 2024b; Kumar et al., 2024; Yao et al., 2024a; Neel and Chang, 2023; Esmradi et al., 2023) and discuss the defense and mitigation strategies of LLMs. However, they did not explore which stakeholders and artifacts are affected, which may limit the ability to provide actionable insights and a clear scope for the relevant stakeholders. Additionally, Wang et al. (Wang et al., 2024b) were among the first to outline a research agenda for the LLM supply chain. They highlighted the key components in the LLM supply chain, illustrating them through the interactions within MLOps and DevOps. As the topic is still emerging and rapidly evolving, existing works may not comprehensively incorporate insightful ideas and perspectives from online preprints or web blogs. Therefore, we believe that a comprehensive LLM supply chain overview as well as a taxonomy of risks and clear guidance of mitigation in the LLM supply chain could be beneficial for both industrial and academic audiences. It could guide practitioners to pay attention to the LLM supply chain risks and provide a visionary map for academic researchers to devise new mitigation or defensive techniques.
In our paper, we provide a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types, as illustrated in Figure 1. Our definition of the LLM supply chain differs from Want et al.’s work (Wang et al., 2024b) in emphasizing the LLM supply chain more on the basic components and their supply relationships between them. We developed taxonomies of risk types, risky actions, and mitigations related to various LLM supply chain stakeholders and components. Specifically, we fist holistically collected both academic and online resources. Then, we summarize the risk scenarios as stakeholders performing risky actions on specific supply chain components, which in turn lead to distinct risk types. We illustrate the risk scenarios from the perspective of upstream contributors, downstream users, and administrators. Furthermore, we list the types of various mitigation measures in response to the LLM supply chain risks.
@@ 2.2.5. Applications
Applications provide functionalities and services directly used by users. These applications include chatbots, web applications, intelligent agents, text-generation tools, and more. Application developers integrate large models into real-world products to develop intelligent applications with API interfaces. For example, the LLM plugin store can offer a wide range of customizable plugins that allow developers to easily access and implement various functionalities tailored to specific needs. This ecosystem is examined in detail in Zhao et al.’s forward-looking analysis of the LLM app store (Zhao et al., 2024a).
@@ 4.3. Taxonomy of Risky Actions
4.3.1. Risks on Code
The code, including packages, frameworks, and plugins, have become essential at every stage of LLM application development. However, using these libraries comes with potential risks.
Implant Malicious Code. Attackers can develop malicious code and attempt to implant it through various methods, finally integrating it into the LLM supply chain. For instance, Zhao et al. (Zhao et al., 2024b) revealed 9 malicious dataset loading scripts on Hugging Face, and perform root cause analysis of vulnerable formats. It can cause the risk of malicious use (LR2).